# Towards Fair RAG: On the Impact of Fair Ranking in Retrieval-Augmented Generation

### v1

@@ C.2 Normalization of EE-R
The lower bound is achieved when $\mathbf{\epsilon}$ becomes $\mathbf{\epsilon}^{-}$, which is an exposure vector of the worst case deterministic ranking $\sigma^{-}$ (permutations that rank all non-relevant items above relevant items). Given the assumption made from equation 8 and, $\textrm{C}^{+}$ and $\textrm{C}^{-}$, which are set of indices of relevant and non-relevant items, respectively,
@@
Let $\mathbf{\epsilon}^{*}_{d}$ be the exposure of a relevant items, $S^{*}_{n}$ be the set of all optimal rankings, $w_{\sigma^{\prime}}$ be the weight on $\sigma\in S^{*}_{n}$ such that $\sum_{\sigma\in S^{*}_{n}}w_{\sigma^{\prime}}=1$, and $\mathbf{\epsilon}^{\prime}$ be the exposure vector associated with $\sigma^{\prime}$.
@@
$=\mathbf{\epsilon}^{*}_{d}\sum_{i\in\textrm{C}^{+}}\sum_{\sigma^{\prime}\in S^{*}_{n}}w_{\sigma^{\prime}}\mathbf{\epsilon}^{\prime}_{i}$ (17)
$=\mathbf{\epsilon}^{*}_{d}\sum_{\sigma^{\prime}\in S^{*}_{n}}w_{\sigma^{\prime}}\sum_{i\in\textrm{C}^{+}}\mathbf{\epsilon}^{\prime}_{i}$ (18)
$\leq\mathbf{\epsilon}^{*}_{d}\sum_{\sigma^{\prime}\in S^{*}_{n}}w_{\sigma^{\prime}}(m\mathbf{\epsilon}^{*}_{d})$ (since $\mathbf{\epsilon}^{\prime}$ is optimal) (19)
$=\mathbf{\epsilon}^{*}_{d}(m\mathbf{\epsilon}^{*}_{d})$ (since $\sum_{\sigma\in S^{*}_{n}}w_{\sigma^{\prime}}=1$) (20)
@@
With UM, the upper bound of EE-R can be calculated by equation 6.
@@ C.3 Normalization of EU
Recall that the string utility $u_{i}=\mu_{u}(y,\mathcal{G}(\phi_{p}(x,\sigma_{i})))$ is a utility of a RAG model with a sampled ranking $\sigma_{i}\in\Sigma$. Let $u^{(\alpha)}_{i}$ denote the $i$’th string utility when the sampling temperature parameter is set to $\alpha$. Also, let $u^{*}_{i}$ denote the $i$’th string utility, given a ranking $\sigma^{*}_{i}$ that belongs to $\Sigma^{*}$, which is a set of sampled permutations (rankings) from a stochastic oracle retriever. This stochastic oracle retriever is a policy that always places relevant items above non-relevant items, thus having $m!(n-m)!$ number of unique optimal permutations.
To obtain an approximated upper bound of the utility $u_{max}$, when the runs of the experiments were run with $\alpha=(1,2,4,8)$, we take the maximum of $u^{(\alpha)}_{i}$ and $u^{*}_{i}$ for $\forall\alpha$:
$u_{max}=max(u^{(1)}_{1},\cdots,u^{(1)}_{N},u^{(2)}_{1},\cdots,u^{(2)}_{N},u^{(4)}_{1},\cdots,u^{(4)}_{N},u^{(8)}_{1},\cdots,u^{(8)}_{N},u^{*}_{1},\cdots,u^{*}_{N})$ (25)
With $u_{max}$, we max-normalize the EU. Since $\frac{1}{N}\sum^{N}_{i=1}\frac{u_{i}}{u_{max}}$ is the same as $(\frac{1}{N}\sum^{N}_{i=1}u_{i})/u_{max}$, per query $q$, we get a normalized EU
@@ Appendix D Effect of $\alpha$ on Metrics
Figure 6: Effect of a temperature parameter $\alpha$ on the (a) disparity of rankings, (b) ranking quality, and (c) generation quality of RAG. The left column is the results on the LaMP task 1 and the right column is on the LaMP task 4, each corresponding to a classification and text generation task, respectively. RAG model is configured with Contriever and Flan-T5-Base for all six figures. Each data point represents the value of the associated metric for one query.
@@ F.4 Quantification of Fairness-Ranking Quality Tradeoff
Table 3: Values on the left are the gradient of a linear line fit to the data points where x-axis is $\overline{\text{EE-D}}$ and y-axis is $\overline{\text{EE-R}}$. Higher the value, stronger the tradeoff between fairness and ranking quality. Values on the right are the DR-AUC on the disparity-ranking quality ($\overline{\text{EE-D}}$ Vs. $\overline{\text{EE-R}}$) curve. Higher the value, stronger the ranking quality, given consistent tradeoff between fairness and relevance. Oracle retriever is omitted here as $\overline{\text{EE-R}}$ of oracle is almost always 1.0.
@@ G.2 Quantification of the Relationship Between Ranking Quality and Utility
Table 4: Values on the left are the gradient of a linear line fit to the data points where x-axis is $\overline{\text{EE-R}}$ and y-axis is $\overline{\text{EU}}$. Higher the value, stronger the tradeoff between retrieval quality and generation quality. Values on the right are the RU-AUC on the ranking quality-utility ($\overline{\text{EE-R}}$s. $\overline{\text{EU}}$) curve. Higher the value, stronger the general end-performance of a RAG model when every level of relevance is considered. Oracle retriever is omitted here as $\overline{\text{EE-R}}$ of oracle is almost always 1.0.

### v2

@@ C.2 Normalization of EE-R
The lower bound is achieved when $\mathbf{\epsilon}$ becomes $\mathbf{\epsilon}^{-}$, which is an exposure vector of the worst case ranking $\pi^{-}$ (permutations that rank all non-relevant items above relevant items). Given the assumption made from equation 8 and, $\textrm{C}^{+}$ and $\textrm{C}^{-}$, which are set of indices of relevant and non-relevant items, respectively,
@@
Let $\mathbf{\epsilon}^{*}_{d}$ be the exposure of a relevant items, $S^{*}_{n}$ be the set of all optimal rankings, $w_{\pi^{\prime}}$ be the weight on $\pi\in S^{*}_{n}$ such that $\sum_{\pi\in S^{*}_{n}}w_{\pi^{\prime}}=1$, and $\mathbf{\epsilon}^{\prime}$ be the exposure vector associated with $\pi^{\prime}$.
@@
$=\mathbf{\epsilon}^{*}_{d}\sum_{i\in\textrm{C}^{+}}\sum_{\pi^{\prime}\in S^{*}_{n}}w_{\pi^{\prime}}\mathbf{\epsilon}^{\prime}_{i}$ (17)
$=\mathbf{\epsilon}^{*}_{d}\sum_{\pi^{\prime}\in S^{*}_{n}}w_{\pi^{\prime}}\sum_{i\in\textrm{C}^{+}}\mathbf{\epsilon}^{\prime}_{i}$ (18)
$\leq\mathbf{\epsilon}^{*}_{d}\sum_{\pi^{\prime}\in S^{*}_{n}}w_{\pi^{\prime}}(m\mathbf{\epsilon}^{*}_{d})$ (since $\mathbf{\epsilon}^{\prime}$ is optimal) (19)
$=\mathbf{\epsilon}^{*}_{d}(m\mathbf{\epsilon}^{*}_{d})$ (since $\sum_{\pi\in S^{*}_{n}}w_{\pi^{\prime}}=1$) (20)
@@
With MU, the upper bound of EE-R can be calculated by equation 6.
@@ C.3 Normalization of EU
Recall that the string utility $u_{\pi}=\mu_{u}(y,\mathcal{G}(\phi_{p}(x,\pi)))$ is a utility of a RAG model with a sampled ranking $\pi\in\sigma$. Let $\sigma_{\alpha}$ denote a sample of rankings with the fairness control parameter set to $\alpha$. Also, let $\sigma^{*}$ denote the set of sampled permutations (rankings) from the oracle stochastic retriever, a policy that always places relevant items above non-relevant items; thus the oracle generates $m!(n-m)!$ number of unique optimal permutations.
To obtain an approximated upper bound of the utility $u_{max}$, when the runs of the experiments were run with $\alpha=(1,2,4,8)$, we take the maximum over all samples,
$u_{max}=max\left(\{u_{\pi}\}_{\pi\in\sigma_{1}}\cup\{u_{\pi}\}_{\pi\in\sigma_{2}}\cup\{u_{\pi}\}_{\pi\in\sigma_{4}}\cup\{u_{\pi}\}_{\pi\in\sigma_{8}}\cup\{u_{\pi}\}_{\pi\in\sigma^{*}}\right)$ (25)
With $u_{max}$, we max-normalize the EU. Since $\frac{1}{N}\sum_{\pi\in\sigma}\frac{u_{\pi}}{u_{max}}$ is the same as $(\frac{1}{N}\sum_{\pi\in\sigma}u_{\pi})/u_{max}$, per query $q$, we get a normalized EU
@@ Appendix D Effect of $\alpha$ on Metrics
Figure 6: Effect of a fairness control parameter $\alpha$ on the (a) disparity of rankings, (b) ranking quality, and (c) generation quality of RAG. The left column is the results on the LaMP task 1 and the right column is on the LaMP task 4, each corresponding to a classification and text generation task, respectively. RAG model is configured with Contriever and Flan-T5-Base for all six figures. Each data point represents the value of the associated metric for one query.
@@ F.4 Quantification of Fairness-Ranking Quality Tradeoff
Table 3: Values on the left are the gradient of a linear line fit to the data points where x-axis is $\overline{\text{EE-D}}$ and y-axis is $\overline{\text{EE-R}}$. Higher the value, stronger the tradeoff between fairness and ranking quality. Values on the right are the DR-AUC on the disparity-ranking quality ($\overline{\text{EE-D}}$ Vs. $\overline{\text{EE-R}}$) curve. Higher the value, stronger the ranking quality, given consistent tradeoff between fairness and relevance.
@@ G.2 Quantification of the Relationship Between Ranking Quality and Utility
Table 4: Values on the left are the gradient of a linear line fit to the data points where x-axis is $\overline{\text{EE-R}}$ and y-axis is $\overline{\text{EU}}$. Higher the value, stronger the tradeoff between retrieval quality and generation quality. Values on the right are the RU-AUC on the ranking quality-utility ($\overline{\text{EE-R}}$s. $\overline{\text{EU}}$) curve. Higher the value, stronger the general end-performance of a RAG model when every level of relevance is considered.